# The Non Parametric Reconstruction of Binary Black Hole Mass Evolution from GWTC-4.0 Gravitational Wave Catalog

### v1

@@ Abstract
The distribution of binary black hole (BBH) masses and its evolution with redshift provide key insights into the different formation channels of the compact objects and their evolution with cosmic time and stellar properties such stellar metallicity and star formation rate history. We present a non parametric, model-independent joint reconstruction of the redshift evolution of BBH mass distribution from gravitational wave (GW) catalog GWTC-4.0 from the fourth observation of LIGO-Virgo-KAGRA (LVK). This method simultaneously searches for the signature of any linear and quadratic redshift evolution with respect to the low redshift in a Bayesian framework taking into accounting the detector selection effects. We find tentative evidence for a linear redshift-dependent evolution of the mass distribution, consistent over a mass range ($m\gtrsim 50\,M_{\odot}$). While lower mass systems shows no signature of evolution. The quadratic term remains consistent with zero, indicating that a simple linear dependence adequately describes the population up to redshift $z\sim 1$. In future with more GW sources, this technique can shed light into the true nature of the redshift dependence and possibility to uncover subtle evolutionary features in BBH populations and to probe the cosmic history of black hole formation.
1 Introduction
The detection of gravitational waves (GWs) has revolutionized astrophysics by providing a direct means to observe and characterize binary black hole populations across vast cosmic distances. These unprecedented observations have transformed our understanding of how stellar mass black holes form, evolve, and merge over the history of the universe (Abbott et al., 2019, 2016a; Bouffanais et al., 2021; Franciolini et al., 2022; Cheng et al., 2023; Antonelli et al., 2023; Wong et al., 2023; Tiwari & Fairhurst, 2021; Tiwari, 2022, 2025; Kimball et al., 2020, 2021; Afroz & Mukherjee, 2025a, b, c). With each new detection, we gain invaluable information about the masses, spins, and rates of these elusive objects, enabling tests of astrophysical theory in regimes previously inaccessible (Abbott et al., 2021, 2023; LIGO-VIRGO-KAGRA Collaboration, 2025a).
@@ 1 Introduction
Understanding how these astrophysical processes imprint themselves on the black hole mass function as a function of redshift holds the promise of uncovering the origins and evolution of compact object populations, offering insights into star formation history, galaxy evolution, and the end stages of stellar lives. Yet, measuring the redshift dependent mass distribution presents formidable challenges (Spera & Mapelli, 2017; Belczynski et al., 2016; Tanikawa, 2024). Previous studies have explored various approaches to characterize the BBH mass distribution and its potential evolution, ranging from parametric models with specific functional forms to more flexible non-parametric methods (Wysocki et al., 2019; Mandel et al., 2019; Tiwari & Fairhurst, 2021; Talbot et al., 2019; Wong et al., 2020; Callister & Farr, 2024; Riley & Mandel, 2023; Farah et al., 2025; Talbot et al., 2025; Ray et al., 2023; Magaña Hernandez & Palmese, 2025; Antonini et al., 2025a; Edelman et al., 2023; Heinzel et al., 2025b, 2023, a; Alvarez-Lopez et al., 2025).
Figure 1: Schematic illustration of the impact of selection effects and intrinsic mass evolution on the observable black hole mass distribution. The red regions denote parameter space inaccessible to detectors due to selection bias. The green area represents the detectable mass range under scenarios with no intrinsic evolution, while the yellow region highlights where signatures of positive mass evolution would appear. For negative evolution, part of the green region below the blue line and above the lower red boundary remains observable. Arrows and shaded bands indicate the characteristic observational trends: constant distribution for no evolution, downward shifts for negative evolution, and upward shifts for positive evolution, emphasizing how evolutionary effects manifest relative to the standard detection boundaries.
@@
In order to infer the true physical effect from the data, one needs to take into account the redshift and mass dependent selection effects. Gravitational wave detectors have sensitivity biases favoring the detection of more massive and nearby sources, creating a mass and redshift dependent selection function effect, which can be modeled given the detectors sensitivity (Essick, 2023; Essick & Holz, 2024; Messenger & Veitch, 2013; Gerosa & Bellotti, 2024).
Figure 1 presents a schematic illustration of the black hole mass distribution as a function of redshift, highlighting the role of GW selection effects and evolutionary signatures. The red-shaded regions indicate areas excluded from detection due to sensitivity limitations, while the yellow region represents the detectable parameter space. In the absence of intrinsic evolution, the observed black hole masses should follow a band bounded by the selection threshold and a constant upper mass limit (dashed blue line). However, the presence of mass evolution introduces systematic shifts, positive evolution leads to an increasing upper mass boundary with redshift (dashed black line), while negative evolution reduces it (dashed blue line). The figure thus encapsulates how black hole mass distributions that evolve with cosmic time can be disentangled from selection bias, providing direct observational probes of black hole formation and stellar evolution across the universe.
@@
In this work, we propose a novel, fully non parametric framework for inferring the evolution of the stellar mass black hole mass function using GW observations. Leveraging a Taylor series expansion of the joint mass redshift distribution, this approach makes minimal assumptions about the functional form of the mass function or its evolution. Instead, it allows the data itself to reveal evolutionary patterns, capturing both linear and higher order changes in the population over redshift. This model-independent strategy offers greater flexibility and interpretability, with each Taylor coefficient corresponding to physically meaningful rates and accelerations of population change. Such a framework paves the way for revealing subtle and complex evolutionary signatures that could otherwise remain hidden. Applying this framework to the GWTC-3 (Abbott et al., 2023) and GWTC-4.0 catalog (LIGO-VIRGO-KAGRA Collaboration, 2025b), the most comprehensive collection of BBH mergers to date, we find tentative evidence for mass dependent evolution in the BBH mass function. High mass systems ($40$-$60\,M_{\odot}$) show positive tentative redshift evolution, while lighter black holes exhibit little or no statistically significant evolution. These results are consistent with metallicity driven formation channels and demonstrate the potential of model independent methods for probing black hole populations with future GW observations.
2 Non Parametric Bayesian Inference of the BBH Mass Distribution with Redshift
@@
The aim of this work is to test whether the distribution of BBH component masses evolves with redshift, a question with direct implications for stellar evolution, metallicity, and the astrophysical environments of BBH formation. Gravitational wave observations provide unique access to this distribution, but detector sensitivity introduces strong selection effects: massive systems are detectable at greater distances than lighter ones, biasing the observed sample. Correcting for these biases is essential for reconstructing the intrinsic population. We demonstrate in these sections our Bayesian method in reconstruction of BBH mass distribution from data in a model-independent way.
2.1 Non parametric Mass Redshift Distribution
@@
Rather than relying on parametric models with fixed functional forms, which risk obscuring unanticipated features, we adopt a non parametric, model free framework based on a Taylor expansion of the mass redshift distribution. Instead of prescribing exact shapes, this method expresses $p(m,z)$ as a smooth series expansion in redshift around a reference redshift $z_{\rm ref}$:
$p(m,z)=\sum_{n=0}^{N}\frac{1}{n!}\left.\frac{\partial^{n}p(m,z)}{\partial z^{n}}\right|_{z=z_{\rm ref}}(z-z_{\rm ref})^{n},$ (1)
@@
where the truncation order $N$ dictates the level of evolutionary behavior captured, and we choose $z_{\rm ref}=0.25$ to center the expansion of the mass distribution. In this work, we restrict our analysis up to second order term in the Taylor series expansion, justified by the fact that the redshift range of current GW observations extends up to $z\sim 1.0$, along with only a few detected events at high redshift. The resulting approximation is:
$p(m,z)\approx p_{0}(m)+p_{1}(m)(z-z_{\rm ref})+p_{2}(m)(z-z_{\rm ref})^{2},$ (2)
with the coefficient functions defined as
$p_{0}(m)\equiv p(m,z_{\rm ref}),\\ p_{1}(m)\equiv\left.\frac{\partial p}{\partial z}\right|_{z_{\rm ref}},\\ p_{2}(m)\equiv\frac{1}{2}\left.\frac{\partial^{2}p}{\partial z^{2}}\right|_{z_{\rm ref}}.$ (3)
@@
Here, $p_{0}(m)$ encodes the mass distribution at the reference redshift, while $p_{1}(m)$ and $p_{2}(m)$ capture the linear and quadratic evolutionary trends respectively. This construction avoids restrictive parametric assumptions, permits non-monotonic evolution, and provides coefficients with direct physical interpretations.
@@
The recovered evolution coefficients encode distinct aspects of mass function evolution. The linear evolution coefficient $p_{1}(m)$ indicates whether black holes of mass $m$ were more common (positive values) or less common (negative values) at higher redshifts. The magnitude $|p_{1}(m)|$ quantifies the strength of evolution, with $|p_{1}(m)|\sim 1$ corresponding to order-unity changes over $\Delta z\sim 1$. The quadratic evolution coefficient $p_{2}(m)$ captures curvature in evolutionary trends. Positive values of $p_{2}(m)$ indicate accelerating evolution, while negative values suggest deceleration. This term allows for non-monotonic behavior and can reveal whether evolutionary processes are intensifying or saturating with redshift.
2.2 Data Processing and Selection Function
@@
here $\Omega_{m}$ is the matter density and $H_{0}$ is the Hubble constant. We convert each distance sample $d_{L,s}^{(i)}$ to a redshift $z_{s}^{(i)}$ using this relation, mapping the posteriors into the $(m,z)$ plane. We work consistently in the source frame mass distribution for easily understanding the signature of any redshift evolution in the source frame mass distribution.
To calibrate the selection function, we use injection campaigns in which artificial signals of known parameters are added to the detector noise and recovered with the search pipelines. For each injection, we record its source frame masses, luminosity distance, and network signal to noise ratio, classifying it as detected if $\mathrm{SNR}>\rho_{\mathrm{thresh}}$ with $\rho_{\mathrm{thresh}}=10$. We note that while GWTC detections are defined using a false alarm rate threshold, in the injection sets we adopt the SNR cut as the criterion for detectability.
@@
The selection function $S(m,z)$ represents the probability that a binary with component mass $m$ at redshift $z$ will be detected by the GW network. Formally, this is the marginalization over extrinsic parameters $\Omega$ (sky location, orientation, etc.):
$S(m,z)=\int P(\text{det}|m,z,\Omega)\,p(\Omega)\,d\Omega\approx\frac{N_{\mathrm{det}}(m,z)}{N_{\mathrm{inj}}(m,z)},$ (5)
where the integral is estimated empirically from injection campaigns. The LIGO(Abbott et al., 2016b) Virgo(Acernese et al., 2015) KAGRA(Akutsu et al., 2021) (LVK) Collaboration performs extensive injection campaigns by adding simulated GW signals with known parameters to detector noise, then processing these through the same search pipelines used for actual detections. These injections span the parameter space of compact binary coalescences, including various mass ratios, spins, and orientations, providing a comprehensive mapping of detector sensitivity across the $(m,z)$ plane (Abbott et al., 2023; Essick & Holz, 2024; Talbot & Thrane, 2018; Gerosa & Bellotti, 2024; Essick et al., 2025)  [^1] The injection sets and recovered selection functions for GWTC-3 and GWTC-4.0 are publicly available at https://zenodo.org/record/5636816 and https://zenodo.org/records/16740128 respectively.. In practice, we discretize the $(m,z)$ plane into a two dimensional grid of bins $\{(m_{k},z_{j})\}$ with $k=1,\ldots,N_{m}$ and $j=1,\ldots,N_{z}$, computing $S(m_{k},z_{j})=N_{\mathrm{det}}(m_{k},z_{j})/N_{\mathrm{inj}}(m_{k},z_{j})$ where $N_{\mathrm{inj}}$ and $N_{\mathrm{det}}$ are the number of injected and detected signals in each bin. We impose a floor $S(m_{k},z_{j})\geq 0.01$ to avoid numerical instabilities in regions with sparse injections.
2.3 Expected Number of Detections
The key insight is that the observed number of detections in each bin depends on three factors: the intrinsic merger rate density, the mass distribution, and the selection effects. The expected number of detections in bin $(m_{k},z_{j})$ is:

### v2

@@ Abstract
The distribution of binary black hole (BBH) masses and its evolution with redshift provide key insights into the different formation channels of compact objects and their dependence on cosmic time and stellar properties such as metallicity and star formation history. We present a non-parametric framework of the redshift evolution of the BBH mass distribution using the gravitational-wave (GW) catalogs GWTC-3 and GWTC-4 of LIGO-Virgo-KAGRA (LVK). This method simultaneously searches for the signature of any linear and quadratic redshift evolution with respect to the low-redshift population in a Bayesian framework, taking into account the detector selection effects. We find tentative evidence for a linear redshift-dependent evolution of the mass distribution, consistent over a mass range ($m\gtrsim 50\,M_{\odot}$), while lower-mass systems show no significant evolution. The quadratic term remains consistent with zero, indicating that a simple linear dependence adequately describes the population up to redshift $z\sim 1$. With more GW sources in the future, this technique can reveal subtle evolutionary features in BBH populations and provide new insights into the cosmic history of black hole formation.
I Introduction
The detection of gravitational waves (GWs) has revolutionized astrophysics by providing a direct means to observe and characterize binary black hole populations across vast cosmic distances. These unprecedented observations have transformed our understanding of how stellar mass black holes form, evolve, and merge over the history of the universe (Abbott et al., 2019, 2016; Bouffanais et al., 2021; Franciolini et al., 2022; Cheng et al., 2023; Antonelli et al., 2023; Wong et al., 2023; Tiwari & Fairhurst, 2021; Tiwari, 2022, 2025; Kimball et al., 2020, 2021; Afroz & Mukherjee, 2025a, b, c). With each new detection, we gain invaluable information about the masses, spins, and rates of these elusive objects, enabling tests of astrophysical theory in regimes previously inaccessible (Abbott et al., 2021, 2023; LIGO-VIRGO-KAGRA Collaboration, 2025a).
@@ I Introduction
Understanding how these astrophysical processes imprint themselves on the black hole mass function as a function of redshift holds the promise of uncovering the origins and evolution of compact object populations, offering insights into star formation history, galaxy evolution, and the end stages of stellar lives. Previous studies have explored various approaches to characterize the BBH mass distribution and its potential evolution, ranging from parametric models with specific functional forms to more flexible non-parametric methods (Wysocki et al., 2019; Mandel et al., 2019; Tiwari & Fairhurst, 2021; Talbot et al., 2019; Wong et al., 2020; Callister & Farr, 2024; Riley & Mandel, 2023; Farah et al., 2025; Talbot et al., 2025; Ray et al., 2023; Magaña Hernandez & Palmese, 2025; Antonini et al., 2025a; Edelman et al., 2023; Heinzel et al., 2025b, 2023, a; Alvarez-Lopez et al., 2025).
@@
In order to infer the true physical effect from the data, one needs to take into account the redshift and mass dependent selection effects. Gravitational wave detectors have sensitivity biases favoring the detection of more massive and nearby sources, creating a mass and redshift dependent selection function effect, which can be modeled given the detectors sensitivity (Essick, 2023; Essick & Holz, 2024; Messenger & Veitch, 2013; Gerosa & Bellotti, 2024; Essick et al., 2025).
@@
In this work, we propose a fully non parametric framework for inferring the evolution of the stellar mass black hole mass function using GW observations. Leveraging a Taylor series expansion of the joint mass redshift distribution, this approach makes minimal assumptions about the functional form of the mass function or its evolution. Instead, it allows the data itself to reveal evolutionary patterns, capturing both linear and higher order changes in the population over redshift. This strategy offers greater flexibility and interpretability, with each Taylor coefficient corresponding to physically meaningful rates and accelerations of population change. Such a framework paves the way for revealing subtle and complex evolutionary signatures that could otherwise remain hidden. Applying this framework to the GWTC-3 (Abbott et al., 2023) and GWTC-4 catalog (LIGO-VIRGO-KAGRA Collaboration, 2025b), the most comprehensive collection of BBH mergers to date, we find tentative evidence for mass dependent evolution in the BBH mass function. High mass systems ($40$-$60\,M_{\odot}$) very mild indications of positive redshift evolution, while lighter black holes exhibit little or no statistically significant evolution.
These results are consistent with metallicity driven formation channels and demonstrate the potential of non-parametric methods for probing black hole populations with future GW observations.
II Non-Parametric Bayesian Inference of the BBH Mass Distribution with Redshift
@@
The aim of this work is to test whether the distribution of BBH component masses evolves with redshift, a question with direct implications for stellar evolution, metallicity, and the astrophysical environments of BBH formation. Gravitational wave observations provide unique access to this distribution, but detector sensitivity introduces strong selection effects: massive systems are detectable at greater distances than lighter ones, biasing the observed sample. Correcting for these biases is essential for reconstructing the intrinsic population. We demonstrate in this section our Bayesian framework for the reconstruction of the BBH mass distribution from data using a flexible, non-parametric approach.
II.1 Non-Parametric Mass-Redshift Distribution
@@
Rather than relying on parametric models with fixed functional forms, which risk obscuring unanticipated features, we adopt a non-parametric framework based on a Taylor expansion of the mass-redshift distribution. Instead of prescribing exact shapes, this method expresses $\rm{p(m,z)}$ as a smooth series expansion in redshift around a reference redshift $\rm{z_{\rm ref}}$:
$\rm{p(m,z)=\sum_{n=0}^{N}\frac{1}{n!}\left.\frac{\partial^{n}p(m,z)}{\partial z^{n}}\right|_{z=z_{\rm ref}}(z-z_{\rm ref})^{n}},$ (1)
@@
where the truncation order $\rm{N}$ dictates the level of evolutionary behavior captured, and we choose $\rm{z_{ref}=0.25}$ to center the expansion. In this work, we restrict our analysis to the second-order term in the Taylor series expansion, justified by the fact that the redshift range of current GW observations extends up to $\rm{z\sim 1.0}$, with only a few detected events at high redshift. The resulting approximation is:
$\rm{p(m,z)\approx p_{0}(m)+p_{1}(m)(z-z_{\rm ref})+p_{2}(m)(z-z_{\rm ref})^{2}},$ (2)
with the coefficient functions defined as:
$\rm{p_{0}(m)\equiv p(m,z_{\rm ref})},$ $\quad\rm{p_{1}(m)\equiv\left.\frac{\partial p}{\partial z}\right|_{z_{\rm ref}}},$
$\quad\rm{p_{2}(m)\equiv\frac{1}{2}\left.\frac{\partial^{2}p}{\partial z^{2}}\right|_{z_{\rm ref}}}.$ (3)
@@
Here, $\rm{p_{0}(m)}$ encodes the mass distribution at the reference redshift, while $\rm{p_{1}(m)}$ and $\rm{p_{2}(m)}$ capture the linear and quadratic evolutionary trends, respectively. This construction avoids restrictive parametric assumptions, permits non-monotonic evolution, and provides coefficients with direct physical interpretations.
@@
Since the Taylor expansion does not automatically satisfy probability normalization, in our analysis we explicitly normalize the resulting expression at each redshift by dividing by its integral over mass. This ensures that $\rm{\int p(m\mid z,\boldsymbol{\theta})\,dm=1}$ at each redshift, making $\rm{p(m\mid z,\boldsymbol{\theta})}$ a valid conditional probability distribution for mass given redshift. The recovered evolution coefficients encode distinct aspects of mass function evolution. The linear evolution coefficient $\rm{p_{1}(m)}$ indicates whether black holes of mass $\rm{m}$ were more common (positive values) or less common (negative values) at higher redshifts. The magnitude $\rm{|p_{1}(m)|}$ quantifies the strength of evolution. The quadratic evolution coefficient $\rm{p_{2}(m)}$ captures curvature in evolutionary trends. Positive values of $\rm{p_{2}(m)}$ indicate accelerating evolution, while negative values suggest deceleration. This term allows for non-monotonic behavior and can reveal whether evolutionary processes are intensifying or saturating with redshift.
II.2 Data Processing and Selection Function
@@
where $\Omega_{m}$ is the matter density and $H_{0}$ is the Hubble constant. We convert each distance sample $d_{L,s}^{(i)}$ to a redshift $z_{s}^{(i)}$ using this relation, mapping the posteriors into the $(m,z)$ plane. We work consistently in the source frame for easier interpretation of any redshift evolution signatures.
To compute the selection function, we use the publicly released LVK injection campaigns associated with GWTC-3 and GWTC-4  [^1] The injection sets and recovered selection functions for GWTC-3 and GWTC-4 are publicly available at https://zenodo.org/record/5636816 and https://zenodo.org/records/16740128, respectively. These injection campaigns provide simulated compact binary coalescence signals spanning the relevant parameter space. We classify an injection as detected if its network signal-to-noise ratio satisfies $\mathrm{SNR}>10$. We note that while GWTC detections are defined using a false alarm rate (FAR) threshold, we adopt the SNR cut for the injection sets as this threshold is commonly used in population inference studies and provides a conservative approximation to the FAR-based selection criterion for the bulk of the BBH population. This choice is consistent with previous population analyses and ensures computational tractability while capturing the essential sensitivity characteristics of the detector network.
@@
The selection function $\rm{S(m,z)}$ represents the probability that a binary with component mass $\rm{m}$ at redshift $\rm{z}$ is detected by the GW network. Formally, this corresponds to marginalizing the detection probability over extrinsic parameters $\rm{\Omega}$ (e.g., sky location, inclination, and polarization):
$\rm{S(m,z)=\int P(\mathrm{det}\mid m,z,\Omega)\,p(\Omega)\,d\Omega\;\approx\;\frac{N_{\mathrm{det}}(m,z)}{N_{\mathrm{inj}}(m,z)}},$ (5)
where the integral is estimated empirically using the injection campaigns described above. In practice, $\rm{S(m,z)}$ is computed as the fraction of injected signals that are recovered in each region of the $\rm{(m,z)}$ plane. To avoid numerical instabilities in sparsely populated regions of the injection parameter space, we impose a small floor on the selection function and require $\rm{S(m,z)\geq 10^{-3}}$.
Because the injection sets span realistic distributions of mass ratios, spins, and orientations, the resulting selection function automatically incorporates spin- and geometry-dependent detection efficiencies. By working directly with the component mass posteriors $\rm{(m_{1},m_{2})}$, we effectively marginalize over these secondary parameters without imposing restrictive parametric assumptions about their redshift evolution.
II.3 Hierarchical Inference Framework
We develop a hierarchical Bayesian framework that properly accounts for measurement uncertainties in individual event parameters and detector selection effects. Our approach builds on the standard hierarchical inference methodology established in gravitational wave population studies (Mandel et al., 2019; Fishbach et al., 2017; Loredo, 2004), adapted to our non-parametric, binned representation of the mass-redshift distribution. We discretize the continuous $\rm{(m,z)}$ parameter space into a two-dimensional grid of bins $\rm{\{(m_{k},z_{j})\}}$ with $\rm{k=1,\ldots,N_{m}}$ mass bins and $\rm{j=1,\ldots,N_{z}}$ redshift bins.
This bin approach minimizes assumptions about smoothness or correlations between neighboring regions, allowing the data to reveal localized evolutionary features without imposing artificial constraints. The intrinsic number of mergers in bin $\rm{(m_{k},z_{j})}$ is:
$\rm{N_{\rm int}(m_{k},z_{j}\mid\boldsymbol{\theta}_{kj})}$ $=\rm{T_{\rm obs}\,\mathcal{R}(z_{j})\,p(m_{k}\mid z_{j},\boldsymbol{\theta}_{kj})}\,$
$\times\rm{\frac{dV_{c}}{dz}\bigg|_{z_{j}}\,\frac{1}{1+z_{j}}\,\Delta m_{k}\,\Delta z_{j}},$ (6)
where $\rm{T_{\rm obs}}$ is the observation time, $\rm{\mathcal{R}(z_{j})}$ is the merger rate density, $\rm{dV_{c}/dz}$ is the differential comoving volume, and $\rm{(1+z)^{-1}}$ accounts for time dilation. This represents mergers that actually occur in the universe, independent of detection. The expected number of detections is $\rm{N_{\rm exp}(m_{k},z_{j}\mid\boldsymbol{\theta}_{kj})=N_{\rm int}(m_{k},z_{j}\mid\boldsymbol{\theta}_{kj})\times S(m_{k},z_{j})}$, where the selection function $\rm{S(m_{k},z_{j})}$ accounts for detector sensitivity.
For each event $\rm{i}$ and bin $\rm{(k,j)}$, we compute the posterior weight:
$\rm{W_{ijk}=\frac{1}{N_{s}}\sum_{\begin{subarray}{c}\text{samples }\ell\\ \text{in bin }(k,j)\end{subarray}}\frac{1}{\pi(m_{i,\ell},z_{i,\ell})}},$ (7)
where $\rm{\{m_{i,\ell},z_{i,\ell}\}_{\ell=1}^{N_{s}}}$ are the posterior samples for event $\rm{i}$, and $\rm{\pi(m,z)}$ is the parameter estimation prior. This weight represents the fractional posterior support that event $\rm{i}$ has in bin $\rm{(k,j)}$, events.
The hierarchical likelihood for bin $\rm{(k,j)}$ with parameters $\rm{\boldsymbol{\theta}_{kj}}$ is: